# Residual Diffusion Models for Variable-Rate Joint Source Channel Coding of MIMO CSI

### v1

@@ Abstract
Despite significant advancements in deep learning-based CSI compression, some key limitations remain unaddressed. Current approaches predominantly treat CSI compression as a source coding problem, neglecting transmission errors. In finite block length regimes, separate source and channel coding proves suboptimal, with reconstruction performance deteriorating significantly under challenging channel conditions. While existing autoencoder-based compression schemes can be readily extended to support joint source-channel coding, they struggle to capture complex channel distributions and exhibit poor scalability with increasing parameter count. To overcome these inherent limitations of autoencoder-based approaches, we propose Residual-Diffusion Joint Source-Channel Coding (RD-JSCC), a novel framework that integrates a lightweight autoencoder with a residual diffusion module to iteratively refine CSI reconstruction. Our flexible decoding strategy balances computational efficiency and performance by dynamically switching between low-complexity autoencoder decoding and sophisticated diffusion-based refinement based on channel conditions. Comprehensive simulations demonstrate that RD-JSCC significantly outperforms existing autoencoder-based approaches in challenging wireless environments. Furthermore, RD-JSCC offers several practical features, including a low-latency 2-step diffusion during inference, support for multiple compression rates with a single model, robustness to fixed-bit quantization, and adaptability to imperfect channel estimation.
Index Terms:
Generative models, Diffusion models, Compression, CSI, JSCC, variable-rate compression
@@ I Introduction
Deep learning (DL) has significantly advanced various areas within physical layer communication, including non-linear channel code design [2, 3, 4, 5, 6], neural channel decoding [7, 8, 9, 10, 11, 12], and MIMO channel estimation [13, 14, 15]. This work focuses specifically on the challenge of lossy compression of CSI at the physical layer. Traditional compression techniques, such as compressed sensing [16], do not work well CSI compression due to the lack of inherent sparsity in CSI structures, making deep learning a better alternative. The field of lossy compression using neural networks, commonly referred to as neural lossy compression, has gained considerable attention in applications such as image compression [17, 18, 19] and video compression [20]. More recently, similar methodologies have been leveraged to significantly enhance the efficiency of CSI compression [21], starting with CSINet [13], which achieved substantial improvements over the then state-of-the-art compressed sensing methods by leveraging convolutional neural networks (CNNs). This breakthrough led to a series of subsequent studies that further refined CNN-based compression techniques [22, 23, 24, 25, 26]. Further, similar approaches have been adapted for joint source channel coding (JSCC) of CSI [27].
@@
Recently, neural image compression has witnessed a significant breakthrough in both compression efficiency and reconstruction quality with the adoption of diffusion models [28, 29]. Originally designed to generate novel images based on various conditioning variables, these models were rapidly adapted for image compression. In [30], the authors introduced a diffusion-based compression framework that reconstructs images through a reverse diffusion process conditioned on contextual information, outperforming certain GAN-based methods. Further, extremely low rate compression schemes have been proposed by leveraging the strong image priors of pretrained diffusion models [31]. Likewise, [32] utilized vector quantized generative adversarial networks (VQGANs), demonstrating the effectiveness of robust codebook priors for discrete visual feature representation in image synthesis, thus providing new insights for compression. More recently, [19] applied principal component analysis (PCA) to enable variable-rate compression, while [33] proposed leveraging spatial entropy [34] to learn latent representations at multiple resolutions, facilitating a continuous range of bitrates.
@@
While diffusion models have been studied extensively in image compression space, their potential for CSI compression remains comparatively underexplored. A notable advancement in this direction is the generative diffusion-based CSI compression framework proposed in [35], demonstrating significant improvements over traditional autoencoder-based methods for modeling complex channels. Building upon this promising foundation, our work aims to further enhance diffusion-based CSI compression in two key dimensions. First, we prioritize reconstruction fidelity by balancing output diversity, shifting the model’s objective from generating novel samples towards maximizing reconstruction accuracy. Second, we utilize a nested representation learning approach to simultaneously capture multiple compression levels within a unified latent space, enabling dynamic and graceful adaptation of the compression rate in response to fluctuating bandwidth constraints.
@@
  • We propose a residual-diffusion-based formulation that enhances the CSI reconstruction by initializing the reverse diffusion with a coarse CSI estimate, rather than conventional random Gaussian noise.
  • We propose a flexible two-stage decoding framework that adaptively switches between low-complexity autoencoder and high-fidelity diffusion-based decoders, based on channel conditions (Sec. IV).
  • We validate the efficacy of our proposed method by comparing against state-of-the-art deep learning-based JSCC schemes for CSI compression, using the widely recognized COST2100 outdoor dataset [13] and 3GPP indoor dataset using QuaDRiGa [36] (Sec. V).
  • We perform a systematic comparative study across multiple neural architectures and diverse channel scenarios, revealing that diffusion-based models significantly outperform autoencoder-based methods only when channel complexity is sufficiently high (Sec. VII-A).
  • We adopt a nested representation learning framework to support multiple compression levels within a single model (Sec. VI) and support fixed-bit quantization of encoder-output for efficient storage and transmission (Sec. VII-C).
@@ II System Model and Problem Formulation
In this work, we consider a massive MIMO system operating in frequency division duplex (FDD) mode, where a base station (BS) with $N_{t}$ antennas communicates with a user equipment (UE) equipped with $N_{r}$ antennas. Considering the large-scale nature of MIMO, we assume $N_{t}\gg 1$ and set $N_{r}=1$ for simplicity. The downlink CSI is $H_{d}\in\mathbb{C}^{N_{c}\times N_{t}}$ and the uplink CSI is denoted by $H_{u}\in\mathbb{C}^{N_{c}\times N_{t}}$ in the spatial-frequency domain, where $\mathbb{C}$ denotes the set of complex numbers and $N_{c}$ is the number of subcarriers.
The encoder $f_{\text{enc}}$ at the UE is designed to efficiently compress the high-dimensional channel measurement $H_{d}$ into a fixed-length representation $\mathbf{s}\in\mathbb{C}^{k}$. The compression rate is thus given by
$r=\frac{k}{N_{t}\times N_{c}}.$
The compressed representation can be transmitted on a noisy uplink channel using $k$ subcarriers, while imposing an unit power constraint for the subcarriers $\frac{1}{k}\,\mathbb{E}\bigl{[}\mathbf{s}\,\mathbf{s}^{*}\bigr{]}=1$. The received signal at the BS is processed using maximal ratio combining (MRC).
The compressed representation at the receiver is processed in two stages. First, the decoder $f_{\text{dec}}$ reconstructs an estimate $\hat{H}_{d}$ from the compressed representation $\mathbf{s}$, aiming to minimize distortion relative to the original input $H_{d}$. Next, to further refine the reconstruction, the noisy estimate $\hat{H}_{d}$ is processed by a denoising diffusion model $f_{\text{den}}$, iteratively reducing the distortion using the reverse diffusion. Specifically, we use the residual diffusion formulation [37]. Following standard practice in CSI compression literature, we adopt mean squared error (MSE) as the distortion metric.
The encoder function is defined as $f_{\text{enc}}:H_{d}\mapsto\mathbf{s}$, while the decoder function is given by $f_{\text{dec}}:\mathbf{s}\mapsto\hat{H}_{d}$, and the denoising model operates as $f_{\text{den}}:\hat{H}_{d}\mapsto H^{\prime}_{d}$. Given the encoder parameters $\theta_{\text{enc}}$, the compressed representation $\mathbf{s}$ is obtained as $\mathbf{s}=f_{\text{enc}}(H_{d};\theta_{\text{enc}})$. Similarly, the decoder, parameterized by $\theta_{\text{dec}}$, reconstructs an estimate of the target as $\hat{H}_{d}=f_{\text{dec}}(\mathbf{s};\theta_{\text{dec}})$. Finally, the denoising model, governed by $\theta_{\text{den}}$, further enhances the reconstruction, producing $H^{\prime}_{d}=f_{\text{den}}(\hat{H}_{d};\theta_{\text{den}})$. The complete set of model parameters is thus given by $\theta=(\theta_{\text{enc}},\theta_{\text{dec}},\theta_{\text{den}})$.
@@
$\theta^{*}=\arg\min_{\theta}\mathbb{E}_{p(H_{d},H^{\prime}_{d})}\big{[}d(H^{\prime}_{d},f_{\text{den}}(f_{\text{dec}}(f_{\text{enc}}(H_{d};\theta_{\text{enc}});\theta_{\text{dec}});\theta_{\text{den}}))\big{]}$ (1)
@@ III Deep Learning for CSI compression
III-A Autoencoder-based CSI compression
Given the three-dimensional spatially correlated structure of the CSI matrix, using a convolution-based architecture is an efficient choice for compressing the input. We chose a low-complexity design for the auto-encoder. The encoder is implemented as lightweight CNN layers followed by a fully connected layer. To maintain robustness under changing signal-to-noise ratio (SNR) conditions, we use an SNR-adaptation module that dynamically scales the feature activations based on SNR. The complete architecture of both the encoder and the SNR-adaptation block is shown in Fig. 1.
@@ III-A Autoencoder-based CSI compression
We train the autoencoder in a supervised manner to learn an effective low-complexity compression scheme. The encoder compresses the continuous channel measurement $H_{d}$ into a compact latent vector $\mathbf{s}\in\mathbb{C}^{k}$. The decoder reconstructs the approximate channel measurement $\hat{H}_{d}$. The encoder and decoder are jointly optimized in an end-to-end fashion to minimize the reconstruction loss for CSI. We use the MSE loss to enforce reconstruction fidelity by ensuring that the reconstructed output closely matches the input, given by:
$\mathcal{L}_{\text{MSE}}(H_{d},\hat{H}_{d})=\|H_{d}-\hat{H}_{d}\|^{2},$ (2)
[…]
@@ IV Residual Diffusion for CSI Enhancement
In this section, we introduce a framework for enhancing CSI reconstruction at the receiver using residual diffusion [37]. Unlike conventional generative diffusion-based approaches, the proposed method can be seamlessly integrated with both learning-based and traditional non-learning-based techniques, as the diffusion module is trained as a standalone denoising model, making it more suitable for practical adaptation.
@@
Preprocessing.  To reduce computational complexity, we first transform the complex matrix $H_{d}$ from the spatial-frequency domain to the angular-delay domain by applying a two-dimensional inverse fast Fourier transform (2D IFFT). This transformation exploits the inherent sparsity in the angular-delay domain, supported by established assumptions [45]. Subsequently, we retain only the first $32$ elements along the delay dimension, as the remaining coefficients typically approach zero, yielding compact angular-delay domain representations of $H_{d}$. The original CSI matrices can be reconstructed by appending zero matrices of size $32\times(N_{c}-32)$ and performing a 2D FFT. This preprocessing procedure is widely recognized as an efficient CSI representation technique [13, 46, 25]. The performance evaluations reported in this work are conducted in terms of normalized mean squared error (NMSE), following the standard practice in CSI compression literature.
@@
The receiver first obtains a coarse estimate of the channel using the autoencoder described in Sec. III-A, which we refer to as Stage 1. Unlike certain CSI compression methods that quantize the latent representation into a fixed number of bits using vector quantization, we instead map the latent to a fixed-length continuous vector. This continuous representation allows direct mapping of encoded data to the uplink subcarriers and simulates complex channel impairments such as multi-path within the feedback channel.
@@
In Stage 2, the denoising process leverages the output of the autoencoder to initialize the reverse diffusion procedure with the coarse CSI estimate. Following standard practice in the generative modeling literature, we adopt the U-Net architecture [47] as the backbone for the denoising network. Unlike recent generative diffusion-based CSI reconstruction methods [35], which initialize reverse diffusion from random Gaussian noise, we employ a residual diffusion approach that initializes the reverse diffusion with a coarse estimate provided by the autoencoder head and iteratively refines the CSI estimate, as illustrated in Fig. 3.
@@
While initializing with random noise facilitates the generation of diverse samples from the underlying distribution—an objective well-suited for data synthesis—it is suboptimal for compression and reconstruction tasks, where compressed latent features already provide a strong prior. Residual diffusion addresses this by starting from a coarse channel estimate, leading to more accurate reconstructions. The effectiveness of residual diffusion for reconstruction tasks has been well demonstrated in the context of image compression [48]. We now formally describe the complete CSI compression and reconstruction pipeline.
@@
In Stage 1, the encoder produces a compressed representation $\mathbf{s}$, which is then used to obtain a coarse reconstruction $\hat{H}_{d}$ of the input channel $H_{d}$. During the denoising stage, the objective is to further refine $\hat{H}_{d}$ by sampling $\mathbf{Z}\sim p(\mathbf{z}\mid\hat{H}_{d})$. This leads to the formulation of a residual denoising diffusion process, given by
$p(\mathbf{z}_{0:T}\mid\hat{H}_{d})=p(\mathbf{z}_{T})\prod_{t=1}^{T}\mathcal{N}(\mathbf{z}_{t-1};\mu_{\theta}(\mathbf{z}_{t},\hat{H}_{d},t),\beta_{t}\mathbf{I}),$ (3)
@@
where $\mathbf{z}_{0:T}=(\mathbf{z}_{0},\dots,\mathbf{z}_{T})$ denotes a realization of the stochastic process $(\mathbf{Z}_{0},\dots,\mathbf{Z}_{T})$, $\mu_{\theta}$ is the learnable mean function parameterized by $\theta$, $\beta_{t}$ defines the variance schedule at each time step $t$, and $\mathbf{I}$ is the identity matrix.
@@
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t},\quad t=1,2,\dots,T,$ (4)
where $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,I)$, $\alpha_{t}=1-\beta_{t}$, and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. As $t$ increases, the noisy latent variable $\mathbf{z}_{t}$ gradually converges to a standard Gaussian distribution. Typically, $T$ is chosen as a relatively large value (e.g., 20–50), with the reverse diffusion process initialized from pure Gaussian noise. However, in the context of CSI compression, this strategy is suboptimal, as a coarse channel estimate $\hat{H}_{d}$ can be readily obtained using either conventional compression techniques [1] or low-complexity deep learning-based methods [49].
To exploit the availability of the coarse estimate $\hat{H}_{d}$, we adopt a residual diffusion approach as presented in [48], which modifies the initialization step as
$\mathbf{z}_{N}=\sqrt{\bar{\alpha}_{N}}\hat{H}_{d}+\sqrt{1-\bar{\alpha}_{N}}\boldsymbol{\epsilon}_{N},$ (5)
@@
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\left(\mathbf{z}_{0}+\eta_{t}\mathbf{r}\right)+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t},\quad t=1,2,\dots,N,$ (6)
where $\mathbf{r}$ denotes the residual between the clean channel $\mathbf{z}_{0}=H_{d}$ and the coarse estimate $\hat{H}_{d}$, i.e., $\mathbf{r}=\hat{H}_{d}-\mathbf{z}_{0}$. The weighting sequence $\{\eta_{t}\}_{t=1}^{N}$ is designed such that $\eta_{1}\to 0$ and $\eta_{N}=1$.
@@
$\mathbf{z}_{t-1}=k_{t}\mathbf{z}_{0}+m_{t}\mathbf{z}_{t}+\sigma_{t}\boldsymbol{\epsilon},$ (7)
@@
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\left(\mathbf{z}_{0}+\lambda\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}\mathbf{r}\right)+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t},$ (9)
@@
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}(\lambda\mathbf{r}+\boldsymbol{\epsilon}_{t}).$ (10)

### v2

@@ Abstract
Despite significant advancements in deep learning-based CSI compression, some key limitations remain unaddressed. Current approaches predominantly treat CSI compression as a source-coding problem, thereby neglecting transmission errors. Conventional separate source and channel coding suffers from the cliff effect, leading to significant deterioration in reconstruction performance under challenging channel conditions. While existing autoencoder-based compression schemes can be readily extended to support joint source-channel coding, they struggle to capture complex channel distributions and exhibit poor scalability with increasing parameter count. To overcome these inherent limitations of autoencoder-based approaches, we propose Residual-Diffusion Joint Source-Channel Coding (RD-JSCC), a novel framework that integrates a lightweight autoencoder with a residual diffusion module to iteratively refine CSI reconstruction. Our flexible decoding strategy balances computational efficiency and performance by dynamically switching between low-complexity autoencoder decoding and sophisticated diffusion-based refinement based on channel conditions. Comprehensive simulations demonstrate that RD-JSCC significantly outperforms existing autoencoder-based approaches in challenging wireless environments. Furthermore, RD-JSCC offers several practical features, including a low-latency 2-step diffusion during inference, support for multiple compression rates with a single model, robustness to fixed-bit quantization, and adaptability to imperfect channel estimation.
@@ I Introduction
Deep learning (DL) has significantly advanced various areas within physical layer communication, including non-linear channel code design [26, 39, 22, 1, 4], neural channel decoding [41, 43, 12, 17, 3, 16], and MIMO channel estimation [53, 13, 45]. This work focuses specifically on the challenge of lossy compression of CSI at the physical layer. Traditional compression techniques, such as compressed sensing [27], do not work well CSI compression due to the lack of inherent sparsity in CSI structures, making deep learning a better alternative. The field of lossy compression using neural networks, commonly referred to as neural lossy compression, has gained considerable attention in applications such as image compression [5, 6, 31] and video compression [30]. More recently, similar methodologies have been leveraged to significantly enhance the efficiency of CSI compression [15], starting with CSINet [53], which achieved substantial improvements over the then state-of-the-art compressed sensing methods by leveraging convolutional neural networks (CNNs). This breakthrough led to a series of subsequent studies that further refined CNN-based compression techniques [52, 32, 36, 38, 24]. Furthermore, similar approaches have been adapted for joint source-channel coding (JSCC) of CSI [54].
@@
Recently, neural image compression has witnessed a significant breakthrough in both compression efficiency and reconstruction quality with the adoption of diffusion models [18, 46]. Originally designed to generate novel images based on various conditioning variables, these models were rapidly adapted for image compression. In [55], the authors introduced a diffusion-based compression framework that reconstructs images through a reverse diffusion process conditioned on contextual information, outperforming certain GAN-based methods. Furthermore, extremely low-rate compression schemes have been proposed by leveraging the strong image priors of pretrained diffusion models [8]. Likewise, [40] utilized vector quantized generative adversarial networks (VQGANs), demonstrating the effectiveness of robust codebook priors for discrete visual feature representation in image synthesis, thus providing new insights for compression. More recently, [31] applied principal component analysis (PCA) to enable variable-rate compression, while [29] proposed leveraging spatial entropy [9] to learn latent representations at multiple resolutions, facilitating a continuous range of bitrates and [47] proposed a dictionary-based quantization approach with variable length masks to support multiple compression rates.
@@
While diffusion models have been studied extensively in image compression, their potential for CSI compression remains underexplored. A notable advancement in this direction is the generative diffusion-based CSI compression framework proposed in [25], demonstrating significant improvements over traditional autoencoder-based methods for modeling complex channels. Building upon this promising foundation, our work aims to further enhance diffusion-based CSI compression in two key dimensions. First, we prioritize reconstruction fidelity by balancing output diversity, shifting the model’s objective from generating novel samples towards maximizing reconstruction accuracy. Second, we employ a nested representation learning approach to capture multiple compression levels within a unified latent space. By using continuous-valued latents with scalar quantization, our method naturally handles noisy transmission, as it maps directly to subcarriers with power constraints. This design enables dynamic and graceful adaptation of the compression rate in response to fluctuating bandwidth constraints.
@@
  • We propose RD-JSCC, a flexible two-stage framework that initializes reverse diffusion with a coarse autoencoder estimate and iteratively refines the residual error. This design enables adaptive switching between lightweight autoencoder decoding and 2-step diffusion-based refinement to balance reconstruction quality with low-latency inference requirements.(Sec. IV)
  • We adopt a nested representation-learning framework based on continuous-valued latent spaces that supports multiple compression rates within a single unified model, enabling fine-grained bandwidth adaptation without retraining or model-switching overhead (Sec. V).
  • We validate RD-JSCC against state-of-the-art deep learning-based JSCC schemes using the COST2100 outdoor and 3GPP indoor datasets, demonstrating order-of-magnitude performance improvements over non-diffusion baselines and establishing that diffusion-based refinement is particularly effective for statistically complex channel distributions (Sec. VI).
  • We conduct comprehensive ablation studies that quantify the contributions of the residual diffusion formulation, encoder complexity, and SNR adaptation, and systematically analyze the complexity-performance trade-offs across different channel scenarios to determine when diffusion-based refinement is justified (Sec. VII-A).
  • We address practical deployment considerations by demonstrating that NMSE improvements translate to substantial BLER reductions, validating robustness to imperfect channel estimation, and demonstrating fixed-bit quantization down to 4 bits per latent dimension with negligible performance degradation (Sec. VI-F, Sec. VII-B, Sec. VII-C).
@@ II System Model and Problem Formulation
In this work, we consider a massive MIMO system operating in frequency-division duplex (FDD) mode, in which a base station (BS) with $N_{t}$ antennas communicates with user equipment (UE) with $N_{r}$ antennas. Following conventions in existing literature and available datasets, we consider massive MIMO systems where $N_{t}\gg 1$ and set $N_{r}=1$, resulting in a SIMO configuration from each user’s perspective. This single-antenna receiver assumption simplifies CSI feedback while maintaining practical relevance, as the base station can still serve multiple users simultaneously through spatial multiplexing. Our proposed techniques readily extend to full MIMO scenarios with $N_{r}>1$ by incorporating an additional antenna dimension in the deep learning architecture. The downlink CSI is $\bm{H}_{d}\in\mathbb{C}^{N_{c}\times N_{t}}$ and the uplink CSI is denoted by $\bm{H}_{u}\in\mathbb{C}^{N_{c}\times N_{t}}$ in the spatial-frequency domain, where $\mathbb{C}$ denotes the set of complex numbers and $N_{c}$ is the number of subcarriers.
The encoder $f_{\text{enc}}$ at the UE is designed to efficiently compress the high-dimensional channel measurement $\bm{H}_{d}$ into a fixed-length representation $\mathbf{s}\in\mathbb{C}^{k}$. The compression rate is thus given by
$\rho=\frac{k}{N_{t}\times N_{c}}.$
The compressed representation can be transmitted on a noisy uplink channel using $k$ subcarriers, while imposing a unit power constraint for the subcarriers $\frac{1}{k}\,\mathbb{E}\bigl[\mathbf{s}\,\mathbf{s}^{*}\bigr]=1$. The received signal at the BS is processed using maximal ratio combining (MRC).
The compressed representation at the receiver is processed in two stages. First, the decoder $f_{\text{dec}}$ reconstructs an estimate $\hat{\bm{H}}_{d}$ from the compressed representation $\mathbf{s}$, aiming to minimize distortion relative to the original input $\bm{H}_{d}$. Next, to further refine the reconstruction, the noisy estimate $\hat{\bm{H}}_{d}$ is processed by a denoising diffusion model $f_{\text{den}}$, iteratively reducing the distortion using the reverse diffusion. Specifically, we use the residual diffusion formulation [34]. Following standard practice in CSI compression literature, we adopt mean squared error (MSE) as the distortion metric.
The encoder function is defined as $f_{\text{enc}}:\bm{H}_{d}\mapsto\mathbf{s}$, while the decoder function is given by $f_{\text{dec}}:\mathbf{s}\mapsto\hat{\bm{H}}_{d}$, and the denoising model operates as $f_{\text{den}}:\hat{\bm{H}}_{d}\mapsto\bm{H}^{\prime}_{d}$. Given the encoder parameters $\theta_{\text{enc}}$, the compressed representation $\mathbf{s}$ is obtained as $\mathbf{s}=f_{\text{enc}}(\bm{H}_{d};\theta_{\text{enc}})$. Similarly, the decoder, parameterized by $\theta_{\text{dec}}$, reconstructs an estimate of the target as $\hat{\bm{H}}_{d}=f_{\text{dec}}(\mathbf{s};\theta_{\text{dec}})$. Finally, the denoising model, governed by $\theta_{\text{den}}$, further enhances the reconstruction, producing $H^{\prime}_{d}=f_{\text{den}}(\hat{\bm{H}}_{d};\theta_{\text{den}})$. The complete set of model parameters is thus given by $\theta=(\theta_{\text{enc}},\theta_{\text{dec}},\theta_{\text{den}})$.
@@
$\begin{split}\theta^{*}=\arg\min_{\theta}\mathbb{E}_{p(\bm{H}_{d},\bm{H}^{\prime}_{d})}\big[&d\big(\bm{H}_{d},f_{\text{den}}(f_{\text{dec}}(f_{\text{enc}}(\bm{H}_{d};\theta_{\text{enc}});\\ &\theta_{\text{dec}});\theta_{\text{den}})\big)\big]\end{split}$ (1)
@@ III Deep Learning for CSI compression
Given the three-dimensional, spatially correlated structure of the CSI matrix, a convolution-based architecture is an efficient choice for input compression. We chose a low-complexity design for the auto-encoder. The encoder is implemented as lightweight CNN layers followed by a fully connected layer. To maintain robustness under varying signal-to-noise ratio (SNR) conditions, we employ an SNR-adaptation module that dynamically scales feature activations based on SNR. The complete architecture of both the encoder and the SNR-adaptation block is shown in Fig. 1.
We employ continuous latent representations instead of vector quantization (VQ) because VQ suffers from the cliff effect: single-bit errors in codebook indices can cause catastrophic reconstruction failure when transmitted over noisy channels. Continuous representations provide graceful degradation: as channel noise increases, reconstruction errors scale proportionally, thereby supporting joint source-channel coding without requiring additional error protection layers. This approach enables direct modulation onto uplink subcarriers while maintaining compatibility with scalar quantization when needed, as demonstrated in Section VII-C.
@@
We train the autoencoder in a supervised manner to learn an effective low-complexity compression scheme. The encoder compresses the continuous channel measurement $\bm{H}_{d}$ into a compact latent vector $\mathbf{s}\in\mathbb{C}^{k}$. The decoder reconstructs the approximate channel measurement $\hat{\bm{H}}_{d}$. The encoder and decoder are jointly optimized end-to-end to minimize the CSI reconstruction loss. We use the MSE loss to enforce reconstruction fidelity by ensuring that the reconstructed output closely matches the input, given by:
$\mathcal{L}_{\text{MSE}}(\bm{H}_{d},\hat{\bm{H}}_{d})=\|\bm{H}_{d}-\hat{\bm{H}}_{d}\|^{2},$ (2)
where $\|\cdot\|$ denotes the Frobenius norm.
@@ IV Residual Diffusion for CSI Enhancement
In this section, we introduce a framework for enhancing CSI reconstruction at the receiver using residual diffusion [34]. Unlike conventional generative diffusion-based approaches, the proposed method can be seamlessly integrated with both learning-based and non-learning-based techniques, as the diffusion module is trained as a standalone denoising model, thereby facilitating practical adaptation.
@@
Preprocessing. To reduce computational complexity, we first transform the complex matrix $\bm{H}_{d}$ from the spatial-frequency domain to the angular-delay domain by applying a two-dimensional inverse fast Fourier transform (2D IFFT). This transformation exploits the inherent sparsity in the angular-delay domain, supported by established assumptions [50]. Subsequently, we retain only the first $32$ elements along the delay dimension, as the remaining coefficients typically approach zero, yielding compact angular-delay domain representations of $\bm{H}_{d}$. The original CSI matrices can be reconstructed by appending zero matrices of size $32\times(N_{c}-32)$ and performing a 2D FFT. This preprocessing procedure is widely recognized as an efficient technique for CSI representation [53, 51, 38]. The performance evaluations reported in this work are conducted in terms of normalized mean squared error (NMSE), following the standard practice in CSI compression literature.
@@
The receiver first obtains a coarse estimate of the channel using the autoencoder described in Sec. III, which we refer to as Stage 1. Unlike CSI compression methods that use vector quantization to discretize the latent representation into fixed-bit codewords, we map the channel directly to a fixed-length, continuous-valued latent vector. This continuous representation enables direct mapping of each latent dimension to uplink Orthogonal Frequency-Division Multiplexing (OFDM) subcarriers for feedback transmission from the UE to the base station, and naturally captures complex channel impairments, such as multipath fading, within the feedback channel.
@@
In Stage 2, the denoising process leverages the autoencoder output to initialize the reverse diffusion procedure with the coarse CSI estimate. We adopt the U-Net architecture [42] as the backbone of the denoising network, following standard practice in generative modeling. Unlike recent generative diffusion-based CSI reconstruction methods [25] that initialize from random Gaussian noise, we employ a residual diffusion approach. This method initializes the reverse diffusion with the autoencoder’s coarse estimate and iteratively refines the CSI, as illustrated in Figure 3. The decoder adaptively switches between 1-stage and 2-stage decoding based on the uplink channel quality. When the uplink SNR exceeds a predetermined threshold, indicating favorable channel conditions, the decoder terminates at Stage 1. Otherwise, Stage 2 diffusion refinement is invoked to further enhance reconstruction quality.
@@
While initializing with random noise facilitates the generation of diverse samples from the underlying distribution, an objective well-suited for data synthesis, it is suboptimal for compression and reconstruction tasks, where compressed latent features already provide a strong prior. Residual diffusion addresses this by starting from a coarse channel estimate, leading to more accurate reconstructions. The effectiveness of residual diffusion for reconstruction tasks has been well demonstrated in the context of image compression [33]. We now formally describe the complete CSI compression and reconstruction pipeline.
@@
In Stage 1, the encoder produces a compressed representation $\mathbf{s}$, which is then used to obtain a coarse reconstruction $\hat{\bm{H}}_{d}$ of the input channel $\bm{H}_{d}$. During the denoising stage, the objective is to further refine $\hat{\bm{H}}_{d}$ by sampling $\mathbf{Z}\sim p(\mathbf{z}\mid\hat{\bm{H}}_{d})$. This leads to the formulation of a residual denoising diffusion process, given by
$p(\mathbf{z}_{0:T}\mid\hat{\bm{h}}_{d})=p(\mathbf{z}_{T})\prod_{t=1}^{T}\mathcal{N}(\mathbf{z}_{t-1};\mu_{\theta}(\mathbf{z}_{t},\hat{\bm{h}}_{d},t),\beta_{t}\mathbf{I}),$ (3)
@@
where $\hat{\bm{h}}_{d}$ is the current sample of the coarse estimate of the channel. $\mathbf{z}_{0:T}=(\mathbf{z}_{0},\dots,\mathbf{z}_{T})$ denotes a realization of the stochastic process $(\mathbf{Z}_{0},\dots,\mathbf{Z}_{T})$, $\mu_{\theta}$ is the learnable mean function parameterized by $\theta$, $\beta_{t}$ defines the variance schedule at each time step $t$, and $\mathbf{I}$ is the identity matrix.
@@
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{t},\quad t=1,2,\dots,T,$ (4)
where $\bm{\epsilon}_{t}\sim\mathcal{N}(0,I)$, $\alpha_{t}=1-\beta_{t}$, and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. As $t$ increases, the noisy latent variable $\mathbf{z}_{t}$ gradually converges to a standard Gaussian distribution. Typically, $T$ is chosen as a relatively large value (e.g., 20–50), with the reverse diffusion process initialized from pure Gaussian noise. However, in the context of CSI compression, this strategy is suboptimal, as a coarse channel estimate $\hat{\bm{H}}_{d}$ can be readily obtained using either conventional compression techniques [44] or low-complexity deep learning-based methods [48].
To exploit the availability of the coarse estimate $\hat{\bm{H}}_{d}$, we adopt a residual diffusion approach as presented in [33], which modifies the initialization step as
$\mathbf{z}_{N}=\sqrt{\bar{\alpha}_{N}}\hat{\bm{h}}_{d}+\sqrt{1-\bar{\alpha}_{N}}\bm{\epsilon}_{N},$ (5)
@@
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\left(\mathbf{z}_{0}+\eta_{t}\mathbf{r}\right)+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{t},\quad t=1,2,\dots,N,$ (6)
where $\mathbf{r}$ denotes the residual between the clean channel $\mathbf{z}_{0}=\bm{h}_{d}$ and the coarse estimate $\hat{\bm{h}}_{d}$, i.e., $\mathbf{r}=\hat{\bm{h}}_{d}-\mathbf{z}_{0}$. The weighting sequence $\{\eta_{t}\}_{t=1}^{N}$ is designed such that $\eta_{1}\to 0$ and $\eta_{N}=1$.
@@
$\mathbf{z}_{t-1}=k_{t}\mathbf{z}_{0}+m_{t}\mathbf{z}_{t}+\sigma_{t}\bm{\epsilon},$ (7)
@@
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\left(\mathbf{z}_{0}+\lambda\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}\mathbf{r}\right)+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{t},$ (9)
@@
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}(\lambda\mathbf{r}+\bm{\epsilon}_{t}).$ (10)